If the rise is
3/2 and the run is 5, then what is
the ratio of the rise to the run?

Answer by Alan3354(69443)

If the rise is
3/2 and the run is 5, then what is
the ratio of the rise to the run?
----------------
It's the rise/run = 1.5/5 = 3/10 or 0.3